OpenStudy (anonymous):

find two numbers whose sum is 40 and whose product is a maximum answer pls

OpenStudy (anonymous):

answer pls

OpenStudy (tbnrfrags):

one sec

OpenStudy (anonymous):

what ?

OpenStudy (tbnrfrags):

both numbers are 20

OpenStudy (anonymous):

can you answer it with solution pls ? i want to see it .. pls :)

OpenStudy (tbnrfrags):

x+y=40 xy=p y=40-x x(40-x)=P 40x-x^2=P P'=40-2x -40=-2x x=20 20+y=40 20=y when x and y = 20 you get the maximum product

OpenStudy (anonymous):

thank you handsome :)

OpenStudy (tbnrfrags):

<3